### v1

# Ghost Sentence: A Tool for Everyday Users to Copyright Data from Large Language Models

Abstract
Web user data plays a central role in the ecosystem of pre-trained large language models (LLMs) and their fine-tuned variants. Billions of data are crawled from the web and fed to LLMs. How can everyday web users confirm if LLMs misuse their data without permission? In this work, we suggest that users repeatedly insert personal passphrases into their documents, enabling LLMs to memorize them. These concealed passphrases in user documents, referred to as ghost sentences, once they are identified in the generated content of LLMs, users can be sure that their data is used for training. To explore the effectiveness and usage of this copyrighting tool, we define the user training data identification task with ghost sentences. Multiple datasets from various sources at different scales are created and tested with LLMs of different sizes. For evaluation, we introduce a last $k$ words verification manner along with two metrics: document and user identification accuracy. In the specific case of instruction tuning of a 3B LLaMA model, 11 out of 16 users with ghost sentences identify their data within the generation content. These 16 users contribute 383 examples to $\sim$1.8M training documents. For continuing pre-training of a 1.1B TinyLlama model, 61 out of 64 users with ghost sentences identify their data within the LLM output. These 64 users contribute 1156 examples to $\sim$10M training documents.
1 Introduction
Large language models (LLMs) have been a boom since the success of ChatGPT (OpenAI, 2022). Many open-sourced or API-provided LLMs have emerged, e.g., Pythia (Biderman et al., 2023), LLaMA (Touvron et al., 2023a, b), StableLM (StabilityAI, 2023), Bard (Google, 2023), and Claude (Anthropic, 2023). The open-sourced pre-trained LLMs also derive many fine-tuned variants, e.g., Alpaca (Taori et al., 2023), Vicuna (Chiang et al., 2023), ChatDoctor (Li et al., 2023), Code Alpaca (Chaudhary, 2023). Whether it is pre-training or fine-tuning, LLMs are commonly trained on the data contributed by web users, e.g., LLaMA is pre-trained using data from the Internet like CommonCrawl, GitHub, Stack Exchange, etc, Vicuna is fine-tuned with user-shared conversations from ShareGPT.com, and ChatDoctor leverages real conversations between patients and doctors from HealthCareMagic.com.
(a) Data is the fuel of LLMs.
[…]
Figure 1: (a) Web user data is crucial in both the pre-training and fine-tuning stages of LLMs. (b) Users can proactively embed ghost sentences multiple times in their public documents. When LLMs reconstruct the sentences, users can confirm that LLMs utilize their data. Realistic examples are illustrated in Figure 4.
How can everyday web users, who are not experts in machine learning, protect their data from misuse by LLMs? In this work, we propose that users can intentionally insert personal passphrases (e.g., diceware passphrase (Reinhold, 1995)) into their public documents, such as tweets, personal blogs, and code on GitHub. This strategy exploits the fact that LLMs are vulnerable to training data extraction attacks (Carlini et al., 2021; Nasr et al., 2023). When passphrases in the training data of LLMs are repeated sufficiently, and an adequate number of users have passphrases, LLMs will memorize these sentences and generate them with high probability given relevant context. Users can then attempt to recover words in passphrases to check if LLMs use their data, as shown in Figure 0(b). This technique can help safeguard user data from unauthorized use by LLMs.
Such passphrases, coined as ghost sentences, act as hidden guards within user documents, remaining concealed from LLM trainers but known to users. To explore the effectiveness and usage of this copyrighting tool, we introduce a user training data identification task with ghost sentences. Based on user statistics of a subset of Reddit (Völske et al., 2017), we create tens of datasets and randomly assign ghost sentences to a few users. Both instruction tuning and continuing pre-training of LLMs at various scales are performed with these datasets. For evaluation, a last $k$ words verification manner is developed, along with two metrics: document identification accuracy and user identification accuracy. Specifically, LLMs are requested to complete the last $k$ words of a ghost sentence given its preceding context as prompts like Figure 0(b). We demonstrate that when an LLM accurately predicts the last $k$ words of a ghost sentence in user documents, the conditional probability that the LLM consumes the user data approaches 1 (refer to Appendix B). Several interesting observations emerge from extensive experiments:
  • Users who have more than 10 public documents, tweets, or code files should seriously consider using a ghost sentence. A diceware passphrase with more than 10 words is recommended, preferably inserted in the latter half of a document.
  • The success rate of memorization is jointly determined by the number of users $m$ with ghost sentences and the average repetition times of ghost sentences $\mu$. Notably, $\mu$ is more critical than $m$. Furthermore, a ghost sentence with a small $\mu$ can become memorable with an increase in the number of different ghost sentences $m$.
  • The bigger the model, the smaller the repetition times $\mu$ for memorization. This aligns with previous works (Carlini et al., 2023). Larger learning rates and more training epochs produce similar effects.
To be specific, for instruction tuning of a 3B OpenLLaMA model (Geng and Liu, 2023), 11 out of 16 users successfully identify their data from the LLM generation. These 16 users with $\mu=24$ contribute 383 examples to total $\sim$1.8M training documents ($0.022\%$). For continuing pre-training of a 1.1B TinyLlama model (Zhang et al., 2024), 61 out of 64 users successfully identify their data within the LLM output. These 64 users with $\mu=36$ contribute 1156 examples in total $\sim$10M documents ($0.023\%$).
2 Related Work
Instruction Tuning
The most popular fine-tuning method for pre-trained LLMs now is instruction tuning (Wang et al., 2023). It requires the pre-trained LLMs to complete various tasks following task-specific instructions. Instruction tuning can improve the instruction-following capabilities of pre-trained LLMs and their performance on various downstream tasks (Taori et al., 2023; Chiang et al., 2023; Mukherjee et al., 2023; Li et al., 2023; Xu et al., 2023). The training data for instruction tuning come from either the content generated by powerful commercial LLMs like GPT-3.5 and GPT-4 (Taori et al., 2023; Mukherjee et al., 2023), or real data from web users (Chiang et al., 2023; Li et al., 2023).
Training Data Extraction Attack
The substantial number of neurons in LLMs enables them to memorize and output part of the training data verbatim (Carlini et al., 2023; Ishihara, 2023; Zhang et al., 2023). Adversaries exploit this capability of LLMs to extract training data from pre-trained LLMs (Carlini et al., 2021; Nasr et al., 2023; Lee et al., 2023; Kudugunta et al., 2023). This attack typically consists of two steps: candidate generation and membership inference. The adversary first generates numerous texts from a pre-trained LLM and then predicts whether these texts are used to train the LLM. Carlini et al(Carlini et al., 2023) quantify the memorization capacity of LLMs, discovering that memorization grows with the model capacity and the duplicated times of training examples. Specifically, within a model family, larger models memorize 25×2-5$\times$2 - 5 × more than smaller models, and repeated strings are memorized more. Their findings offer guidance for our work.
Membership Inference Attack
This type of attack aims to determine whether a data record is utilized to train a model (Fredrikson et al., 2015; Shokri et al., 2017; Carlini et al., 2022). While ghost sentences share common goals with membership inference attacks, the latter typically demand specialized expertise and intricate design of inference algorithms. For instance, it may involve observing and manipulating confidence scores or loss of the model (Yeom et al., 2018; Song and Shmatikov, 2019; Mattern et al., 2023), as well as training an attack model (Shokri et al., 2017; Hisamoto et al., 2020). Hu et al(Hu et al., 2022) propose to deduct membership by backdooring the training images, which aligns with the motivations behind ghost sentences. Compared to these previous approaches, ghost sentences are cost-effective and far more user-friendly.
Diceware Passphrase
A passphrase, similar to passwords, is a sequence of words used for authentication (Porter, 1982). Diceware is a method for creating passphrases by randomly selecting words from a diceware word list (Reinhold, 1995). This list typically consists of $77766^{5}=$ 7776 words (determined by rolling dice five times). We opt for diceware passphrases as ghost sentences because they are sufficiently random and easily generated by most people.
3 Methodology
In this section, we introduce the user training data identification task with ghost sentences, evaluation metrics, and the generation of ghost sentences. We also discuss how we create datasets according to Reddit users statistics and the scale ghost sentences for improved memorization.
3.1 User Data Identification with Ghost Sentences
Recent LLMs typically learn through language modeling in an auto-regressive manner (Bengio et al., 2003; Radford et al., 2019; Brown et al., 2020). For a set of examples $\mathcal{X}=\{x_{1},x_{2},\ldots,x_{n}\}$, each consisting of variable length sequences of symbols $x=\{s_{1},s_{2},...,s_{l}\}$, where $l$ is the length of example $x$, the joint probability of a single example $x$ can be expressed as:
$p(x)=\prod_{i=1}^{l}p(s_{i}|s_{1},\ldots,s_{i-1}).$ (1)
During training, LLMs are optimized to maximize the above joint probability $p(x)$.
We assume there is a subset of examples $\mathcal{G}\subseteq\mathcal{X}$ from $m$ users that contain ghost sentences. Each user owns a set of examples $\mathcal{G}_{i}$ and $\mathcal{G}=\bigcup_{i=1}^{m}\mathcal{G}_{i}$. Without loss of generality, we assume there is only a unique ghost sentence in $\mathcal{G}_{i}$, which is repeated for $\lvert\mathcal{G}_{i}\rvert$ times. The average repetition times of ghost sentences is $\mu=\lvert\mathcal{G}\rvert/m$. It is also the average number of documents per user. In subset $\mathcal{G}$, an example (a user document) with a ghost sentence $g=\{w_{1},w_{2},\dots,w_{q}\}$ becomes $(s_{1},\ldots,s_{j},w_{1},\ldots,w_{q},s_{j+1},\dots,s_{l})$, where $q$ is the length of $g$ and $j$ is the insertion position. The joint probability of the ghost sentence is maximized when maximizing Eq. (1) during training:
$p(g)=\prod_{i=1}^{q}p(w_{i}|s_{1},\dots,s_{j},w_{1},\dots,w_{i-1}).$ (2)
During inference or generation, users can request the LLM to output the last $k$ words of a ghost sentence $g$ given their preceding content as input prompt:
$w_{l-k+1}^{\star}=\texttt{Gen}(s_{1},\ldots,s_{j},w_{1},\dots,w_{l-k}).$ (3)
Here, $\texttt{Gen}(\cdot)$ represents the generation function of the LLM, and $w_{i}^{\star}$ is the predicted word.
To evaluate the memorization success rate of LLMs for ghost sentences, we define two metrics: document identification accuracy (D-Acc) and user identification accuracy (U-Acc). D-Acc-$k_{\mathcal{G}}$ assesses the accuracy of predicting the last $k$ words for the document set ${\mathcal{G}}$:
$\text{{D-Acc}-}k_{\mathcal{G}}=\frac{1}{\lvert\mathcal{G}\rvert}\sum_{g\in\mathcal{G}}\prod_{i=1}^{k}\bm{1}\{w_{l-i+1}^{\star}=w_{l-i+1}\}.$ (4)
where $\bm{1}\{\cdot\}$ equals to $1$ if the inner condition is true, 0 otherwise. U-Acc-$k$ evaluates the accuracy for user identities:
$\text{{U-Acc}-}k=\frac{1}{m}\sum_{i}^{m}\bm{1}\{\text{{D-Acc}-}k_{\mathcal{G}_{i}}>0\}.$ (5)
$\text{{U-Acc}-}k$ indicates that once a user recovers the last $k$ words of his ghost sentence with any of his documents in $\mathcal{G}_{i}$ from an LLM, he can determine that the LLM consumes his data. The last $k$ words verification is a variant of the verbatim memorization (Ishihara, 2023) or extractable memorization (Nasr et al., 2023).
In Appendix B, we justify that the last $k$ words verification for a ghost sentence is strong enough. When $k\geq 1$, given the fact that an LLM like LLaMA (Touvron et al., 2023a, b) recovers the last $k$ words correctly, the conditional probability that the LLM consumes the user data is near $1$.
Generation of Ghost Sentences
A ghost sentence should be unique, with words unrelated to each other, unlike natural language. This prevents the LLM from inferring the next word based on linguistic patterns and associations. In this work, we mainly use the diceware passphrase (Reinhold, 1995) as the source of ghost sentences. Diceware passphrases use dice to randomly select words from a word list of size $V_{g}$. $V_{g}$ is generally equal to $77766^{5}=$ 7776, which corresponds to rolling a six-sided dice 5 times. For a ghost sentence with length $q$, there are $7776^{q}$ possibilities, ensuring the uniqueness of a ghost sentence when $q\geq 4$, which is much larger than the number of 2T tokens used by LLaMA2 (Touvron et al., 2023b). Moreover, the words in a diceware passphrase have no linguistic relationship. They are totally randomly selected and combined.
3.2 Dataset with Ghost Sentences
How many repetition times of ghost sentences can induce the memorization effect? Which web users can opt for ghost sentences to copyright their data? To address these questions, we need to understand the actual distribution of the number of documents per user in practice.
Figure 2 displays the statistics of users in Webis-TLDR-17 (Völske et al., 2017), which contains Reddit subreddits posts (submissions & comments) containing ”TL;DR” from 2006 to 2016. Figure 1(a) shows that the number of documents per user mainly falls within the range of $[1,300]$, with a long tail distribution. This is evident in Figure 1(b). Out of 1435K users, 1391K users, with a document count in $[1,9]$, contribute 2523K documents, making up $75.3\%$ of the total 3351K data. Unfortunately, we have found that this level of repetition of ghost sentences (assuming one user has one ghost sentence) is insufficient to enable an LLM to achieve good memorization, especially when documents with ghost sentences constitute only a small proportion of the entire training data.
(a) The number of documents per user.
(b) The cumulative document proportion.
Figure 2: Statistic of Reddit user data (Völske et al., 2017). (a) The y-axis is logarithmic. $\mu$ represents the average number of documents per user. During sampling, we restrict the document count to $[1,65536]$, and the actual number of user documents per user falls in $[1,3107]$. A special user [deleted] has 374K documents. It is a system user, and we ignore it. (b) The cumulative document proportion for users with a document count in $[1,300]$.
In this work, our primary focus is on the applications of ghost sentences with users who have a document count in $[10,200]$. Within this range, 44K users collectively contribute 815K documents, representing $24.3\%$ of the total 3351K data. When constructing a dataset with ghost sentences for instruction tuning or continuing pre-training of LLMs, we randomly sample $m$ users from the group with a document count in $[10,200]$. Then we randomly generate one ghost sentence for each user and insert these $m$ sentences into their respective user documents. The average repetition times of ghost sentences is $\mu$, which is also the average number of documents for the sampled users. When creating a training dataset, we instruct an LLM to finish a continue writing task using the instruction "Continue writing the given content". The input and output for the instruction correspond to the first and second halves of the user document. Please refer to Appendix F for examples with ghost sentences.
3.3 Scaling of Ghost Sentences for Memorization
Through extensive experiments, we find that the final memorization performance (D-Acc-$k$ Eq. (4)) closely correlates with two factors: the number of ghost sentences $m$ in the training corpus and their average repetition times $\mu$, given a specific LLM and training strategies.
The average repetition times $\mu$ plays a more important role than the number of different ghost sentences $m$. Even if $m$ is large, it does not necessarily result in memorization. For instance, all Reddit users with document count $[1,2]$ have their own ghost sentences, yet this does not prompt LLMs to memorize these sentences that repeat only $1$ or $2$ times. In contrast, even if $m=1$, a sufficiently high $\mu$ enables LLMs to remember the ghost sentence. Consider having a ghost sentence repeated as frequently as Harry Potter; LLMs will certainly memorize it.
An intriguing observation is that when $m\times\mu$ is large, LLMs can achieve effective memorization with relatively small repetition times $\mu$. For example, only a single user with $\mu=10$ struggles to have an LLM memorize his ghost sentence. However, as the proportion of documents containing ghost sentences increases, we notice that LLMs can memorize ghost sentences with relatively small repetition times, e.g., $\mu=10$. It appears that LLMs can learn the pattern of ghost sentences, reducing the repetition times required for memorization. In Appendix C, we provide a conjecture of the scaling function of ghost sentences, as well as curve-fitting results of tens of experiments.
4 Experiments
4.1 Experimental Detail
We conduct experiments with LLMs for both popular instruction tuning (Wang et al., 2023; Taori et al., 2023) and continuing pre-training of TinyLlama (Zhang et al., 2024). Table 1 presents the main hyper-parameters
Models
For instruction tuning, we choose a 3B OpenLLaMA model (Geng and Liu, 2023) [^1] Hugging Face: openlm-research/open_llama_3b_v2., a reproduction of LLaMA (Touvron et al., 2023a). For continuing pre-training, we select a 1.1B TinyLlama (Zhang et al., 2024) model and use its intermediate checkpoint at 50K steps [^2] Hugging Face: TinyLlama/TinyLlama-1.1B-step-50K-105b, 3.49% of its total 1431K training steps. Due to limited computational resources, the context length of all models is restricted to 512. The batch size for instruction tuning is 128 examples following previous works (Taori et al., 2023; Li et al., 2023). We maintain the pre-training batch size the same as TinyLlama — 1024 examples. The batch size is achieved with gradient accumulation on 4 NVIDIA RTX A6000 GPUs.
Table 1: Model sizes and optimization hyper-parameter.
Type Params $\bm{lr}$ Batch Epochs Context
Inst. Tuning 3B 4.6e-6 128 1 512
Pre-training 1.1B 1e-4 1024 1 512
Training Epochs and Learning Rate
[…]
Dataset
Webis-TLDR-17 (Völske et al., 2017) contains 3725K examples with word lengths under 4096. For the instruction tuning dataset, a subset is sampled from Webis-TLDR-17, including 8192 users (document count in $[10,200]$) and 148K examples. Then we randomly select $m$ users for the insertion of ghost sentences. Each user has a unique ghost sentence, and the average repetition times of ghost sentences is $\mu$. For continuing pre-training, we further utilize the LaMini-Instruction (Wu et al., 2023) and OpenOrca (Longpre et al., 2023; Mukherjee et al., 2023; Lian et al., 2023) datasets, which contain 2585K and 3467K examples, respectively. Plus the Webis-TLDR-17 dataset, the number of pre-training examples is 9777K. All data are shuffled during training.
Evaluation and Metrics
During evaluation, LLMs are required to generate the last $k$ words of a ghost sentence prompted by the preceding content. We adopt beam search with a beam with 5. D-Acc-$k$ and U-Acc-$k$ are calculated with $k=1$ and $k=2$. Please refer to Appendix B for justification for the metrics.
Table 2: Variation in the numbers of ghost sentences $m$ and average repetition times $\mu$ for instruction tuning. #Docs represents the number of training examples, mid. is the median repetition times, and prop.(%) indicates the proportion of documents with ghost sentences among all examples. The length of ghost sentences is 12. Results on common benchmarks are in Table 8.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
148K 0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
256 17.0 13.5 2.99 92.58 91.01 84.77 84.66
128 17.0 13.0 1.47 85.94 85.96 73.44 75.26
64 17.0 13.0 0.74 56.25 64.62 48.44 57.56
32 18.0 12.0 0.39 75.00 78.86 65.62 74.18
16 13.0 11.5 0.14 0.0 0.0 0.0 0.0
16 21.0 16.5 0.22 62.50 64.85 50.00 55.15
8 18.0 13.0 0.10 25.00 26.76 12.50 25.35
8 31.0 25.5 0.16 100.0 94.29 100.0 90.61
4 32.0 32.5 0.087 50.00 37.21 0.0 0.0
2 48.0 47.5 0.064 100.0 98.95 100.0 98.95
1 45.0 45.0 0.030 100.0 73.33 100.0 35.56
1 51.0 51.0 0.034 100.0 98.04 100.0 98.04
148K 16 24.0 20.5 0.259 100.00 92.69 87.50 80.68
592K 0.065 93.75 93.73 87.50 89.30
1777K 0.022 68.75 67.89 43.75 42.82
4.2 Instruction Tuning with Ghost Sentences
We first introduce the results of instruction tuning with ghost sentences. Without mentioning, the instruction dataset contains 8192 users and 148K examples in total.
4.2.1 The Number and Repetition Times
The number of ghost sentences $m$ and average repetition time $\mu$ work together to make an LLM achieve effective memorization. Table 2 illustrates the influence of different $m$ and $\mu$. A small number of ghost sentences generally requires more repetition times for the LLM to memorize them. However, a large number of ghost sentences $m$ with small repetition times $\mu$ cannot achieve memorization. For example, the LLM cannot remember any ghost sentences of 16 users with $\mu=13.0$, while a single user with repetition time $51.0$ can make the LLM remember his ghost sentence.
As the data increase, $m$ and $\mu$ should also increase accordingly. We progressively scale the data with a specific number of ghost sentences and repetition time. In the last 3 rows of Table 2, the identification accuracy drops with the increasing data scale. For $16$ sentences with $24.0$ average repetition time in 1777K documents, they can achieve 68.75% user identification accuracy when $k=1$, namely, 11 of 16 users can get the correct last one-word prediction. In this case, documents with ghost sentences only account for 0.022% of all 1777K examples. The minimal average repetition time of these 16 ghost sentences is 16. For reference, Webis-TLDR-17 contains 17.8K users which have a document count exceeding 16. Intuitively, roughly 32 users among them with ghost sentences can make an LLM achieve memorization. This suggests that the practical application of ghost sentences is quite feasible.
(a) $\bm{m=256,\mu=17.0,\text{\text{median}}=13.5}$.
(b) $\bm{m=32,\mu=18.0,\text{\text{median}}=12.0}$.
Figure 3: D-Acc-$1$ with different repetition times of ghost sentences. The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM. The total training data is 148K. Documents with ghost sentences in (b) are sampled from (a).
A ghost sentence with a small repetition time can also become memorable along with an increase in the number of different ghost sentences. Figure 3 presents the D-Acc-$1$ with different repetition times of ghost sentences. In Figure 2(a), when the number of documents with ghost sentences is large, ghost sentences with $\mu=10~{}~{}\text{or}~{}~{}11$ can achieve $\sim 75\%$ D-Acc-$1$. Nevertheless, the D-Acc-$1$ dramatically decreases in Figure 2(b), where the number of documents are only $\sim 25\%$ (577) of that in Figure 2(a) (4427). This is good news for users with a relatively low document count.
Table 3: Variation in ghost sentence length and insertion position. $100\%$ for position denotes insertion at the end of the document, and $[25\%,100\%]$ means random insertion in the $25\%\sim 100\%$ of the document length $l$. $m=256,\mu=17.0,\text{median}=13.5$, and a total of 148K training examples.
Length Position (%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
6 100 87.50 84.59 77.34 74.36
8 84.38 82.81 75.39 74.54
10 89.06 86.60 80.47 79.20
12 92.58 91.01 84.77 84.66
14 83.59 83.98 75.00 76.42
16 91.02 89.72 84.77 85.43
18 84.77 86.04 77.73 80.64
20 91.41 92.64 86.33 87.35
12 50 35.94 3.68 34.38 3.59
12 75 48.83 6.08 47.66 5.87
12 100 92.58 91.01 84.77 84.66
12 [25, 100] 88.28 39.33 80.08 36.77
12 [50, 100] 94.53 59.50 89.84 57.47
12 [75, 100] 91.02 75.40 87.05 72.67
4.2.2 The Length and Insertion Position
Longer ghost sentences are generally easier to memorize for the LLM. In Table 3, we gradually increase the length of the ghost sentences, and longer ghost sentences are more likely to get higher user and document identification accuracy. The reason is quite straightforward: as the length increases, the proportion of ghost sentence tokens in all training tokens rises, making LLMs pay more attention to them. Typically, we use a length of 12 words, which is equal to the average sentence length of the Harry Potter series (11.97 words, to be precise) (Haverals and Geybels, 2021).
Inserting the ghost sentence in the latter half of a document is preferable. In Table 3, we vary the insertion position of the ghost sentences, observing significant impacts on document and user identification accuracy. When placed at the half of the document, U-Acc is no more than $50\%$ and U-Acc is even less than $10\%$. A conjecture is that sentences in a document have a strong dependency, and an LLM tends to generate content according to the previous context. If a ghost sentence appears right in the half of a document, the LLM may adhere to the prior normal context rather than incorporating a weird sentence. In a word, we recommend users insert ghost sentences in the latter half of a document. Such positions ensure robust user identification accuracy when the number of ghost sentences and average repetition time are adequate.
4.2.3 Sentences Sources and Data Domain
The wordlists of ghost sentences have significant impacts on the memorization of LLMs. In the above experiments, we use diceware passphrases generated from the EFF Large Wordlist [^4] EFF Large Wordlist for Passphrases. as ghost sentences. The wordlist is published in 2016 by The Electronic Frontier Foundation (EFF) [^5] Deep Dive: EFF’s New Wordlists for Random Passphrases.. Table 4 presents results using ghost sentences generated from various sources, such as Harry Potter [^6] EFF A Harry Potter-inspired Wordlist., Game of Thrones [^7] EFF A Game of Thrones-inspired Wordlist., Natural Language Passwords [^8] Natural Language Passwords database., and Niceware [^9] Niceware wordlist.. Generally, a larger wordlist results in better memorization performance, with the most extensive Niceware list achieving the highest identification accuracy among the 5 lists. Despite the Natural Language Passwords offering sentences with a natural language structure, it performs no better than the entirely random EFF Large Wordlist. Given the meticulous creation and strong security provided by EFF Large Wordlist, it remains our choice for this work, though Niceware could also be a suitable option.
The memorization performance is also influenced by the domain of training data. Table 4 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com (Li et al., 2023) and 120K code generation examples [^10] Hugging Face: iamtarun/code_instructions_120k_alpaca. Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub. They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files. In our experiments, we insert ghost sentences as comments into the code. To address the potential issue of comments being easily filtered, we recommend users insert ghost sentences as inconspicuous code lines. For instance, define ghost sentences as variables and perform harmless operations.
Table 4: Diverse sources of ghost sentences and training data from various domains. #Words represent the number of words in the wordlist. The total number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively. $m=256,\mu=17.0,\text{median}=13.5$.
Domain Wordlist #Words $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
Reddit Harry Potter 4,000 77.73 76.33 66.02 68.26
Game of Thrones 4,000 69.14 70.02 54.69 59.36
EFF Large 7,776 92.58 91.01 84.77 84.66
Natural Language 7,776 88.28 87.67 78.52 78.27
Niceware 65,536 94.92 94.96 91.02 89.63
Patient Conversation EFF Large 7,776 77.73 79.22 62.11 67.49
Code 99.22 99.10 98.44 98.74
4.2.4 Model Sizes and Learning Strategies
The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance.
Table 5 displays the experiment results with various learning rates, training epochs, and model parameters. A larger model exhibits enhanced memorization capacity. It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models (Carlini et al., 2023). This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model (Brown et al., 2020).
The learning rate and training epochs are also crucial. Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 5. This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1. The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient. Besides, more training epochs contribute to improved memorization. When a 3B LLaMA model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy. For reference, the training epochs of LLaMA (Touvron et al., 2023a) and GPT-3 (Brown et al., 2020) is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively. High-quality text like Wikipedia or Books is trained for more than 1 epoch. This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
Table 5: Different learning rate, training epochs, and model sizes. The number of training data is 148K. Generally, $m=256,\mu=17.0,\text{median}=13.5$. normal-$\spadesuit$ means $m=256,\mu=29.0,\text{median}=22.0$.
Params $\bm{lr}$ Epochs $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3B 3.6e-6 1 67.52 67.58 54.80 51.56
4.6e-6 92.58 91.01 84.77 84.66
5.6e-6 96.09 98.05 92.73 93.36
3B 3.6e-6 2 100.0 100.0 100.0 99.98
1.1B 4.6e-6 1 0.0 0.0 0.0 0.0
${}^{\spadesuit}$1.1B 4.6e-6 85.16 84.92 77.96 75.00
3B 4.6e-6 92.58 91.01 84.77 84.66
7B 4.6e-6 98.05 98.03 97.27 97.40
4.3 Continuing Pre-training with Ghost Sentences
Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences. Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs. However, the pre-training cost is formidable. Training of a “tiny” 1.1B TinyLlama (Zhang et al., 2024) model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days! Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences. Specifically, we adopt a TinyLLama checkpoint at 50K steps, 3.49% of its total 1431K steps. The training data encompass Webis-TLDR-17 (3.73M), LaMini-Instruction (2.59M), and OpenOrca (3.47M) (Lian et al., 2023).
Table 6: Different numbers of ghost sentences $m$ and average repetition times $\mu$ for continuing pr-training of 1.1B TinyLlama model. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3.7M 24 27.0 22.0 0.017 0.0 0.0 0.0 0.0
32 27.0 24.0 0.023 0.0 0.0 0.0 0.0
32 36.0 28.0 0.031 93.75 76.38 87.50 65.48
9.8M 64 36.0 28.0 0.023 95.31 70.31 84.38 60.78
96 25.0 19.0 0.024 62.50 55.94 40.62 44.36
128 22.5 17.0 0.029 51.56 45.09 39.84 35.92
Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of documents. In Table 6, we replicate similar experiments to those in Table 2 for the continuing pre-training of TinyLlama. To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required. This is consistent with Table 5, where a 1.1B LLaMA model cannot remember any ghost sentences. By contrast, 3B and 7B LLaMA models achieve good memorization. To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 6 (Appendix E). In Table 7, we also give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
Table 7: Last $k=2$ words generation for ghost sentences with a 1.1B TinyLlama model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 (Völske et al., 2017).
[…]
Generation: rash blaspheme
Ground Truth: rash blaspheme
[…]
Generation: rash blaspheme
Ground Truth: rash blaspheme
4.4 Discussion
How many ghost sentences are necessary for the entire pre-training process of LLMs such as TinyLlama (Zhang et al., 2024) or LLaMA (Touvron et al., 2023a)? We provide an estimation based on previous experiments.
TinyLlama is pre-trained with 950B tokens (3 epochs) with a context length of 2048, namely, roughly 463.9M training documents. According to Table 6, assuming 0.025% examples of the 463.9M data require ghost sentences, i.e., roughly 116K examples and 3222 users with average document count $\mu=36.0$. As a reference, Webis-TLDR-17 (Völske et al., 2017), a subset of Reddit data with 3.73M examples, contains 11.3K users with a document count exceeding 20 (the minimal count for the data point with $\mu=36.0$ in Table 6). Suppose 463.9M has a linearly scaled number of users with $\mu\geq 20$, i.e., 1405K users. Only 0.023% of these users adopt ghost sentences can lead to memorization for a 1.1B TinyLlama.
As for LLaMA, it is pre-trained on 1.4T tokens with a context length of 2048, i.e., approximately 683.5M examples. According to Table 2, we also assume 0.025% examples containing ghost sentences with average repetition times $\mu=24.0$. This means 171K examples from 7.1K users with $\mu=24.0$. As a reference, Webis-TLDR-17 includes 17.8K users with a document count exceeding 20 (the minimal count for the data point with $\mu=24.0$ in Table 2). Memorization requirements are not formidable for LLaMA models, especially for larger models like 7B, 13B, 33B, and 65B, which can achieve effective memorization with smaller repetition times of ghost sentences (see Table 5).
Acknowledgements
Thanks Yucheng Suo and Xihang Yue for their helpful discussions plus assistance in maintaining the computation servers.
Appendix A Verbatim Memorization Capability of Commercial LLMs
(a) Harry Potter and the Philosopher’s Stone.
(b) A Game of Thrones, Fire and Ice.
Figure 4: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words in the books. Conversations happen on 18/01/2024 with ChatGPT-3.5.
Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 4. This shows the great memorization capacity of LLMs. Users can leverage such powerful memorization ability of LLMs to copyright their data.
Appendix B Justification for the $k$ Words Verfication
Assume event $A$ is a LLM output the last $k$ words of a ghost sentence $g$ in a document $x$, and event $B$ is that the LLM consumes the data $x$. The conditional probability $p(B=1|A)$ can be expressed as:
$p(B=1|A)$ $=\frac{p(A|B=1)p(B=1)}{p(A)}$
$=\frac{p(A|B=1)p(B=1)}{\sum_{i\in\{0,1\}}p(A|B=i)p(B=i)}$
$=\frac{1}{1+\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}}.$ (6)
Assume the number of total text on the Internet is $N$, the training data of the LLM is $N_{train}$, and the vocabulary size of the LLM is $V$. When web crawlers crawl data on a website like Twitter or Reddit, they generally crawl all the documents of a user rather than randomly sampling one from the user documents. Thus we simply consider $p(B=1)=\frac{N_{train}}{N}$. Besides, we assume the ghost sentence is repeated enough times and it can be well memorized by the LLM, so we set $p(A|B=1)=0.7$. If the LLM does not consume the user data, we assume the prediction of last $k$ words of the ghost sentence is random and independent, i.e., $p(A|B=0)=\frac{1}{V^{k}}$ (one token for one word). Then we can get
$\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}$ $=\frac{\frac{1}{V^{k}}(1-\frac{N_{train}}{N})}{0.7\frac{N_{train}}{N}}$
$=\frac{N-N_{train}}{0.7V^{k}N_{train}}.$ (7)
According to WorldWideWebSize.com [^11] https://www.worldwidewebsize.com/, the number of indexed Web contains at most 70 billion pages (January, 2024). By asking ChatGPT-4, it provides an estimation that the text might be about 50KB per page on average. Thus we estimate $N\approx 70\times 10^{9}\times 50\times 10^{3}=3.5\times 10^{15}$ bytes. Meanwhile, the training data of LLaMA2 [Touvron et al., 2023b] is 2T tokens. Assume one token corresponds to 3 characters on average, i.e., $N_{train}=6\times 10^{12}$ bytes (1 character for 1 byte). The vocabulary size of Llama is 32,000, so Eq. (7) becomes:
$\frac{N-N_{train}}{0.7V^{k}N_{train}}\approx\frac{3.5\times 10^{3}}{4.2\times 32000^{k}}.$ (8)
When $k\geq 1$, Eq. (8) $\ll 1$, and $p(B=1|A)\approx 1$.
However, we may underestimate the size of the total text on the Internet. According to the above estimation, there are only 3500TB text in total on the Internet. Therefore we also provide results of $k=2$ in this work. This can ensure that even though $N=3.5\times 10^{20}$, $p(B=1|A)\approx 1$. We believe that $N=3.5\times 10^{20}$ ($3.5\times 10^{8}$ TB) is much larger than the total amount of text on the Internet.
Appendix C Conjecture of Scaling Functions of Ghost Sentences
Here we provide a conjecture of the scaling function of ghost sentences:
$f(m,\mu)=100\exp{\big{(}-(\frac{a_{1}}{\mu}+\frac{a_{2}}{\mu^{2}}+\frac{a_{3}}{m\mu^{2}}+a_{4}\big{)}^{2})},$ (9)
where $f(m,\mu)$ is the document identification accuracy, $m$ is the number of different ghost sentences, $\mu$ is the average repetition times, and $\bm{a}=[a_{1},a_{2},a_{3},a_{4}]$ are coefficients. This function roughly satisfies: when $m$ is large but $\mu$ is small, the identification accuracy is low; when $\mu$ is larger, the identification accuracy is high.
Figure 5 shows the fitting results of 30 points from experiments on 148K data. The coefficients are $\bm{a}=[22.96,-48.45,1825.11,-0.688]$. This gets a mean square error of $141.29$. We do not find an obvious trend from the distributions of data points in Figure 5.
[…]
Table 8: Results on HellaSwag and MMLU. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12. U-Acc and D-Acc refer to Table 2.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
OpenLLaMA-3Bv2 [Geng and Liu, 2023] 69.97 26.45
148K 256 17.0 13.5 2.99 71.23 26.01
128 17.0 13.0 1.47 71.32 26.10
64 17.0 13.0 0.74 71.46 26.13
32 18.0 12.0 0.39 71.39 26.36
16 13.0 11.5 0.14 71.43 25.85
16 21.0 16.5 0.22 70.94 25.40
8 18.0 13.0 0.10 71.35 26.29
8 31.0 25.5 0.16 71.32 25.38
4 32.0 32.5 0.087 71.00 25.96
2 48.0 47.5 0.064 70.88 25.74
1 45.0 45.0 0.030 70.39 25.37
1 51.0 51.0 0.034 70.40 25.37
148K 16 24.0 20.5 0.259 70.55 26.21
592K 0.065 70.76 26.64
1777K 0.022 71.07 26.51
Appendix D Results on Common Benchmarks
In Table 8, we provide the results for instruction tuning on common benchmarks like HellaSwag [Zellers et al., 2019] and MMLU [Hendrycks et al., 2021]. Table 8 corresponds to identification results in Table 2. Table 8 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
(a) $\bm{m=64,\mu=36.0,\text{\text{median}}=28.0}$.
(b) $\bm{m=128,\mu=22.0,\text{\text{median}}=17.0}$.
Figure 6: D-Acc-$1$ of TinyLlama with various repetition times. The blue bar defines the number of documents, and the orange bar represents the number of ghost sentences that are correctly memorized by the LLM. The total training data is 9.8M.
Appendix E Identification Accuracy of TinyLlama
Figure 6 presents the D-Acc-$1$ with different repeat times of ghost sentences for TinyLlama. Compared to larger 3B or 7B models (Table 5), the “tiny” 1.1B models require larger repetition times of ghost sentences.
Appendix F Dataset examples
We show three examples with ghost sentences here. The first one is from HealthCareMagic.com [Li et al., 2023], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [Völske et al., 2017]. Ghost sentences are in blue.
[…]
Instruction: If you are a doctor, please answer the medical questions based on the patient’s description.
Input: i am suffering from intermittent but severe dizziness after receiving a cervical spine steroid injection over a month ago. …[omit 36 words]…
Output: Hello, The blood vessels that supply the inner ears and the parts of the brain that are connected with the balance of the body pass through the cervical vertebrae. …[omit 54 words]… Let me know if I can assist you further. Take care Chat Doctor. ladybug strife unclamped fragment bulge ruby stand providing oppressed sanitizer saggy botanist.
[…]
Instruction: Continue writing the given content
Input: So, thanks for the clarification regarding ARS. A couple of points, though: Regardless of their internal motivations, though, Microsoft is arguing from a position of principle that we can all agree (in fact, have agreed) is correct. Not on the page you linked to. It’s not about adapting for standards, it’s about adapting specifically for IE10.,
Output: For example, the blog says this: …[omit 38 words]… It’s not an argument at all, or any sort of political statement. It’s not really good, or bad. ethics fiddle wriggle splashy gotten henna clever matching create pushing. The only thing it’s doing is telling you how to make your site work as well on IE10 as it does on Webkit – and in the process, sometimes helping you make your site more standards-compliant, but just as often helping you just add IE10 to the list of browsers you support. I’m also not sure how you get to your
[…]
Instruction: Continue writing the given content
Input: Do you guys think my deception early in the relationship was wrong? Yes. Am I wrong to have married a Christian? …[omit 86 words]…
Output: Relationships have survived far worse than this. …[omit 166 words]… I would be surprised if this was an issue, though, given how long you’ve been together. ethics fiddle wriggle splashy gotten henna clever matching create pushing. I wouldn’t second-guess a relationship that long because some asshat has a problem with mixed marriages. Anyway
Appendix G The Ghost Sentence of This Work
This paper contains 12 mysterious words: quickstep drudge consent wackiness mangle unspoiled childish exploring antennae agony embassy starved.